मराठी

If the probability of winning a game is 0.07, what is the probability of losing it?

Advertisements
Advertisements

प्रश्न

If the probability of winning a game is 0.07, what is the probability of losing it?

बेरीज
Advertisements

उत्तर

P(winning the game) = 0.07

P(losing the game) = 1 – 0.07

= 0.93

shaalaa.com
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 16: Probability - VERY SHORT ANSWER TYPE QUESTIONS (VSAQS) [पृष्ठ १६.२९]

APPEARS IN

आर.डी. शर्मा Mathematics [English] Class 10
पाठ 16 Probability
VERY SHORT ANSWER TYPE QUESTIONS (VSAQS) | Q 19. | पृष्ठ १६.२९
Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×